OpenStudy (anonymous):

Using points (0,-5) and (1,-4) find the equation of the line as y =mx + b

OpenStudy (anonymous):

then use point slope equation y-y1=m(x-x1) where you can use x1 and y1 from any of the points and I already told u how to find m-slope

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

m=(y2=y1)/(x2-x1) yep it looks like @Romero is right

OpenStudy (anonymous):

can you confirm my answer y = 1/1x -5

OpenStudy (anonymous):

yea just leave out 1/1

OpenStudy (anonymous):

1/1- 1 and anything times 1 is itself